OpenStudy (anonymous):

A triangular prism has a volume of 5 m3. If each dimension of the prism is tripled, what is the new volume of the prism?

OpenStudy (dumbcow):

V = base*height, where base is a triangle V = 1/2*length*width*height multiply dimensions by 3 V = 1/2*3length*3width*3height 3*3*3 = 27 new volume is 27 times greater